EVA Report Apr 23rd

Crew 245 EVA Report 23APR2021

EVA # 12

Author: Alex Coultrup, Health and Safety Officer

Purpose of EVA: Final survey for emergency shelters and soil samples

Start time: 1015 MDT

End time: 1215 MDT (planned 1300, actual 1200)

Narrative: The EVA this morning started 15 minutes after the scheduled EVA window opened, as the crew had carefully gathered and sterilized required scientific instruments to meet Planetary Protection Protocols. After suiting up, the crew waited 5 minutes before exiting the airlock. The EVA crew departed from the Hab/Hab Grounds by foot, heading northwest. We rapidly identified the ATV road (1103) and followed it north, then west as it curved just before North Ridge, keeping an eye out for potential emergency shelter locations along the way. We followed the trail of GPS points left by the previous day’s EVA crew and demonstrated a successful navigation of our instruments.

After a thirty-minute hike, we reached the top of the western ridge stopping occasionally to drop more points and practice orienting our position with respect to both the GPS device and the map. We took soil samples along the ridge, then admired the view of North Ridge, Schubert’s Pass, and the western buttes. After determining there were no valid candidates for emergency sheltering, the team headed back along the 1103 route pointed out by the outpost as advice. The team managed a very quick return by foot to the Hab and was in the airlock just before 1200.

Destination: Schubert’s Pass

Participants: Health and Safety Officer Coultrup, Crew Scientist Ettlin

Road(s) and routes per MDRS Map: Walking up near a ridge slope by the Hab, taking 1103 walking back.

Mode of travel: Walking.

Green Hab Report – Apr 23rd

Crew 245 GreenHab Report – 23APR2021

GreenHab Officer: Julio A. Hernandez

Environmental control: N/A

Average temperatures: N/A

Hours of supplemental light: 0

Daily water usage for crops: 2 L

Daily water usage for research and/or other purposes: 0

Water in Blue Tank (N/A) gallons

Time(s) of watering for crops: 10am, 4pm

Changes to crops: N/A

Narrative: Greenhab was thoroughly cleaned. Dust on floor and flat surfaces was vacuumed. Crew Botanist, Hernandez, tried to leave the GreenHab in a cleaner state than was originally found.

Harvest: N/A

Support/supplies needed: N/A

Operations Report Apr 23rd

Crew 245 Operations Report 23APR2021

SOL: 13

Name of person filing report: Shravan Hariharan

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Generator: Working nominally

Hours run: 7

From what time last night: 10:00 pm, SOC 62%

To what time this morning: 5:00 am, SOC 100%

List any additional daytime hours when the generator was run: N/A

Solar— SOC 71% at 7:01 pm (Before generator is run at night)

Notes on power system: For the last several sols, the freezers in the Science Dome have been running in order to preserve science samples. Due to this, the station power consumption has been significantly higher than it previously was, resulting in the generator being turned on earlier (as early as 7:00 pm on colder days, and 9:00 pm on warmer days) once the SOC reaches below 70%.

Diesel Reading – 50%

Station Propane Reading – 70%

Water (loft tank): 30 gallons

Water Meter: 150772.1 units

Water (static tank): 25 gallons

Static to Loft Pump used – yes

Water in GreenHab: 0 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: No

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: Yes

Spirit rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: yes, in town

Opportunity rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: no, in town

Curiosity rover used: no

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: yes, in town

Perseverance rover used: No

Hours: 204.6

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: Left tail-light bulb on Perseverance needs to be replaced.

ATV’s Used: None

Reason for use: N/A

Oil Added? No

ATV Fuel Used: 0 gallons

# Hours the ATVs were used today: 0

Notes on ATVs: Mission Director will monitor and refuel ATVs for duration of mission.

HabCar used and why, where?: Nothing to report

CrewCar used and why, where?: Nothing to report

General notes and comments: Nothing to report

Summary of internet: Internet is working nominally. Crew is rationing internet usage during the day to ensure that there is sufficient bandwidth for Mission Support Comms window.

Summary of suits and radios: All radios nominal, all suits currently charging. Suits 9 and 11 used on EVA, and functioned nominally.

Summary of Hab operations: Both decks of Hab were swept and vacuumed. Mopping of lower deck will occur later this evening. EVA suit room was fully cleaned.

Summary of GreenHab operations: Crew Botanist removed experiment/associated equipment from GreenHab, and cleaned/vacuumed entire GreenHab. No further crew activities expected.

Summary of ScienceDome operations: Science Dome was swept, vacuumed, and mopped. Limited soil sample processing remaining for later this evening and tomorrow morning.

Summary of RAM operations: All tools and consumables in RAM were sorted, organized, and put away. Workstations were wiped down and trash was removed. RAM floor was swept, vacuumed, and mopped. No further crew activities expected.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Thank you from all of us in Crew 245 for supporting our mission!

Sol Summary Report Apr 23rd

Crew 245 Sol Summary Report 23APR2021

Sol: 13

Summary Title: Last Day on Mars

Author’s name: Shayna Hume

Mission Status: Executive Officer

Sol Activity Summary:

· 08:30-09:30: Misc. breakfast and briefing

· 09:30-10:00: Prepare for EVA.

· 10:00-12:00: EVA. (ended an hour short) Those in the hab worked on final video recordings and science.

· 12:00-13:30: Lunch!

· 13:30-18:00: Cleaning the hab and wrapping up science!

· 18:00-19:00: Dinner & wrap up reports.

Look Ahead Plan: Tomorrow morning we intend to exit sim soon after we wake up and conduct the check-out process to head to Grand Junction.

Anomalies in work: None.

Weather: Nominal.

Crew Physical Status: No issues.

EVA: EVA #12 report coming in.

Reports to be filed: Operations Report, EVA Report, GreenHab Report

Support Requested: n/a

Copyright © The Mars Society. All rights reserved. | Main Site